Suppose the market demand function for a good is given by.
Q=-5p+1000
i. find the monopolist’s total revenue function
ii. Find maximum total revenue
1
Expert's answer
2020-04-29T09:23:27-0400

Q = -5p + 1000 or p = 200 - 0.2Q.

i. The monopolist’s total revenue function is:

"TR = p*Q = 200Q - 0.2Q^2."

ii. The total revenue is at its maximum if marginal revenue "MR = TR'(Q) = 0."

So:

"200 - 0.4Q = 0,"

"Q = 500" units.